OpenStudy (anonymous):

Find all the zeros of the function f(x)= (x+5)(x-8)^2

OpenStudy (bahrom7893):

just solve this: (x+5)(x-8)^2 = 0

OpenStudy (bahrom7893):

Keep in mind that a product is 0 when one of the multiples is 0.

OpenStudy (anonymous):

x=-5 x=8

OpenStudy (bahrom7893):

so either x+5=0 or (x-8)^2=0

OpenStudy (anonymous):

zeros are basically the value of x

OpenStudy (anonymous):

okay. so I just set the, equal to 0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay thanks